# The Energy Cost of Artificial Intelligence of Things Lifecycle

### v2

@@ VI Energy Cost of Training
In order to train a model, training data and a neural network architecture are needed. The training data typically goes through a preprocessing phase as discussed in Section V and is then split into certain ratios for training and evaluation with a $\beta$ split ratio as:
$N_{\mathrm{S}}=\underbrace{\beta\cdot N_{\mathrm{S}}}_{N_{\mathrm{S,T}}}+\underbrace{(1-\beta)\cdot N_{\mathrm{S}}}_{N_{\mathrm{S,E}}},$ (10)
where $N_{\mathrm{S,T}}$ and $N_{\mathrm{S,E}}$ represent the number of samples for training and evaluation, respectively. While there are many neural network architectures available for developing AIoT models, in this work we focus on four representative types, namely, Multilayer Perceptrons [11], Convolutional Neural Networks [38], Kolmogorov–Arnold Networks [39] and transformers [40]. Training a neural network involves two fundamental phases: forward propagation and backward propagation depicted with blue and red in Fig. 6 on an example MLP. During the forward propagation, input data passes through the network layers. Each layer performs specific tasks, such as computing a linear combination of weights and biases, followed by applying activation functions in dense (fully connected) layers and applying convolutional operations and pooling mechanisms in convolutional layers and pooling layers, respectively. At the end of this phase, the loss is calculated by comparing the predicted outputs to the actual target values. Subsequently, the model performs backward propagation, where it calculates gradients of the loss with respect to each parameter, updates these gradients in reverse through the layers, and adjusts the weights and biases accordingly. This process completes one epoch. Consequently, we can calculate the energy costs associated with forward and backward propagation, leading to an understanding of the overall energy consumption for training, model evaluation, and subsequent inference. In the following sections, we showcase the computational complexity of training and evaluating the aforementioned Neural Networks.
@@ VI-A Computational Complexity of MLP, CNN, KAN and transformer
$M_{\mathrm{MLP}}$ $=\sum_{l=1}^{L-1}\Bigg{(}\underbrace{(2\cdot M_{l-1}\cdot M_{l})}_{\text{Weight and bias}}+\underbrace{2\cdot M_{l}}_{\text{Summation and activation}}\Bigg{)},$ (11)
where the first term in the sum represents one multiplication and one addition ($=2$ FLOPs) related to the weight and bias corresponding to the edges times the number of nodes in that layer ($=M_{l-1}\cdot M_{l}$), while the second term represents the summation of all the values of the incoming edges and the application of the activation function ($=2$ FLOPs) times the number of nodes in that layer ($M_{l}$).
@@
$M_{\mathrm{MLP,FP}}$ $=N_{\mathrm{epochs}}\cdot N_{\mathrm{batch}}\cdot B\cdot M_{\mathrm{MLP}}$ (12)
$=N_{\mathrm{epochs}}\cdot N_{\mathrm{S,T}}\cdot M_{\mathrm{MLP}}.$
@@
$M_{\text{CONV}}=\underbrace{\Bigg{(}\frac{I_{\text{r}}-K_{\text{r}}+2P_{\text{r}}}{S_{\text{r}}}+1\Bigg{)}}_{\text{Output height}}\cdot\underbrace{\Bigg{(}\frac{I_{\text{c}}-K_{\text{c}}+2P_{\text{c}}}{S_{\text{c}}}+1\Bigg{)}}_{\text{Output width}}$ (13)
$\cdot\underbrace{\Bigg{(}C_{in}\cdot K_{\text{r}}\cdot K_{\text{c}}+1\Bigg{)}}_{\text{Computation per filter}}\cdot\underbrace{N_{\text{f}}}_{\text{No. of filters}},$
@@
$M_{\text{POOL}}=\underbrace{\Bigg{(}\frac{I_{\text{r}}-K_{\text{r}}}{S_{\text{r}}}+1\Bigg{)}}_{\text{Output height}}\cdot\underbrace{\Bigg{(}\frac{I_{\text{c}}-K_{\text{c}}}{S_{\text{c}}}+1\Bigg{)}}_{\text{Output width}}\cdot\underbrace{C_{\text{in}}}_{\text{Input channels}}.$ (14)
@@
$M_{\mathrm{CNN,FP}}$ $=N_{\mathrm{epochs}}\cdot N_{\mathrm{S,T}}\cdot$ (15)
@@
$M$ =${}_{\mathrm{KAN}}=\sum_{l=1}^{L-1}(\underbrace{M_{\mathrm{NLF}}\cdot M_{l-1}}_{\text{From B-spline}}$ (16)
$+\underbrace{\left(M_{l-1}\cdot M_{l}\right)}_{\text{Input and output}}\cdot\underbrace{\left[9K\cdot\left(G+1.5K\right)+2G-2.5K+3\right])}_{\text{B-spline and grid}}$
$=\sum_{l=1}^{L-1}(M_{\mathrm{NLF}}\cdot M_{l-1}+\left(M_{l-1}\cdot M_{l}\right)\cdot M_{\mathrm{B}}),$
@@
$M_{\mathrm{KAN,FP}}=$ $N_{\mathrm{epochs}}\cdot N_{\mathrm{S,T}}\cdot M_{\text{KAN}}.$ (17)
@@
Based on the attention mechanism, the transformers [40] represent a paradigm shift in deep learning, replacing traditional architectures such as CNN with self-attention mechanisms. While enabling significant breakthroughs in AI through efficient parallel processing and robust modeling of long-range dependencies, they are relatively complex. In addition to attention layers, the transformer architecture also includes MLP, normalization, and positional encoding, making the analytic expression for the FLOP computation intractable. Following existing literature [43] and approximations  [^3] https://www.gaohongnan.com/playbook/training/how_to_calculate_flops_in_transformer_based_models.html, we approach the estimation as follows: $M_{\text{ATT}}=2\cdot(\underbrace{C\cdot N_{\text{embed}}\cdot 3\cdot N_{\text{embed}}}_{\text{K, Q, V positional embedding}}+\underbrace{C^{2}\cdot N_{\text{embed}}}_{\text{Attention scores}}$ (18) $+\underbrace{N_{\text{head}}\cdot C^{2}\cdot N_{\text{embed}}/N_{\text{head}}}_{\text{reduce}}+\underbrace{C\cdot N_{\text{embed}}\cdot N_{\text{embed}}}_{\text{Projection}}),$ where $C$ is the context length, $N_{\text{embed}}$ is the size of the embedding and $N_{\text{head}}$ is the number of heads. The complexity of the decoder in a transformer is then given by:
$M_{\text{TR}}=N_{\text{decoder\_blocks}}\cdot(M_{\text{ATT}}+\underbrace{4\cdot C\cdot N_{\text{embed}}\cdot FFS}_{\text{MLP blocks}}),$ (19)
@@
$M_{\mathrm{TR,FP}}=$ $N_{\mathrm{epochs}}\cdot N_{\mathrm{S,T}}\cdot M_{\text{TR}},$ (20)
@@ VI-B Model Training
To determine the computational complexity of training a model, also the backward propagation needs to be considered. The standard approximation, also confirmed in [44], is that the backward propagation requires two times the floating-point operations per cycle per core (FLOPs) of the forward propagation. In particular, it requires matrix multiplications for updating the weights and propagating the gradient. Thus, we can approximate the computing complexity of training a model as:
@@ VI-C Model Evaluation
${E_{\mathrm{eval}}=\frac{M_{\mathrm{model}}\cdot N_{\mathrm{S,E}}}{PU_{\mathrm{performance}}}}.$ (23)
@@ VII Energy Cost of Inference
The complexity of making an inference depends on the number of FLOPs required for forward propagating with input sample size $N_{\mathrm{I,P}}$, i.e., $N_{\mathrm{inf}}=M_{\mathrm{model}}\cdot N_{\mathrm{I,P}}$. For example, when we adopt the aforementioned MLP model with $N_{\mathrm{I,P}}=51$, the complexity is calculated as $N_{\mathrm{inf}}=236\cdot 51=12,036$ FLOPs. Therefore, the energy consumption of the forward propagation and the corresponding inference can be calculated as:
$E_{\mathrm{inf}}=\frac{M_{\mathrm{model}}\cdot N_{\mathrm{I,P}}}{PU_{\mathrm{performance}}}.$ (24)
We can then observe from Eq. 23 and 24 that the cost of inference is the same as in evaluation when $N_{\mathrm{S,E}}$ is equal to $N_{\mathrm{I,P}}$. Thus, we can conclude that the computational complexity between finishing one training and inference is on the magnitude of $3\cdot N_{\mathrm{epochs}}\cdot N_{\mathrm{S,T}}/N_{\mathrm{I,P}}$. This indicates that the training process involves significantly more computational operations, particularly when the number of epochs or $N_{\mathrm{S,T}}>>N_{\mathrm{I,P}}$. On the other hand, the energy consumption per bit depends on the model complexity and hardware capability. Moreover, the factor of three in the energy consumption per bit during training versus inference highlights the additional computational load inherent to training.
VIII eCAL: the Energy Cost of AIoT Lifecycle
@@
In this section, we describe the end-to-end energy consumption of an AIoT system for making inferences with the aforementioned neural network architectures. First, we consider the energy consumption of developing the model illustrated on the left in Fig. 1 as follows:
@@
To compare the energy consumption of all the data manipulation components required for developing and deploying an AIoT system, we assume that the models are trained and evaluated using 5-fold cross validation (as discussed in Section VI), they require $256$ samples with input size of $10$, i.e., $N_{\mathrm{S}}\cdot I_{\mathrm{S}}=2560$, collected from an IoT device over a wireless network configured with the following parameters: $P_{\mathrm{T}}=200\leavevmode\nobreak\ [mW]$, $P_{\mathrm{R}}=5\leavevmode\nobreak\ [mW]$, $R_{\mathrm{T}}=R_{\mathrm{R}}=10\leavevmode\nobreak\ [Mbps]$, $N_{\mathrm{\_,cycle},l}=100\leavevmode\nobreak\ [c/b]$, and $P_{\mathrm{\_,cycle}}=10^{-10}\leavevmode\nobreak\ [W/c]$. We also assume the link between the IoT device and the server is stable, and there is no retransmission needed. Additionally, each layer introduces DP and CP overhead of $10\%$ and $5\%$, respectively. Furthermore, the normalization is utilized as data transformation procedure, as discussed in Section V. Finally, model training requires $10$ epochs ($N_{\mathrm{epochs}}=10$) and utilizes $5$-fold validation. The energy consumption of this example is shown in Table III and reveals the contribution of each component to the total energy cost of developing models. Specifically, the data collection dominates the total energy cost for developing a very simple model based on a 3 layer 10 nodes/per layer MLP architecture. For the same parameters, the relative contribution of the data collection decreases for the KAN, CNN and transformer architectures making the energy cost of training the primary contributor. The eCAL calculator enables changing the hyperparameters of the four architectures while also adding well known CNN architectures such as ResNets [45], VGGs [46] or transformer based such as Baichuan 2 [47].
Once the model is developed, it is packaged and deployed in the AIoT system, and it can produce inference in the form of continuous or discrete outputs. During its operational lifecycle, the model will be presented with some input data and requested to produce the corresponding inference. Thus, we can express the associated energy cost as:
@@
This means that we consider the energy cost of inference before the current model needs to be updated. The corresponding energy per bit of inference, once deployed, is calculated as:
@@
Next, to obtain the entire energy cost over the lifetime of a model in the AIoT system, also referred to as $eCAL_{\mathrm{abs}}$ (measured in [$J$]), which is defined as the sum of energy consumed during the model development, $E_{\mathrm{D}}$ and the energy required for performing an inference $E_{\mathrm{inf,p}}$ multiplied by the number of times $\gamma$ the inference is performed with the currently deployed model. In addition, as edge cloud systems may use virtualization in developing and operating the models, we enable considering the impact of such technologies through a factor. As a result, $eCAL_{\mathrm{abs}}$ can be expressed as:
@@
Figure 8: Comparison of energy consumption per bit in different data manipulation components over the lifecycle of the AIoT model.
Figure 9: Energy cost of AIoT lifecycle ($eCAL$) for different models over the number of inferences ($\gamma$), log scale on both axes.
Finally, the proposed metric computing the corresponding energy consumption per bit over the lifecycle of the currently deployed AIoT model can be expressed as:
@@
To further quantify this relationship, we first compare the outcomes of Eq. 26, Eq. 28, and Eq. 30, depicted in Fig. 8, where we assume the virtualization is contributing additional $10\%$ to the total energy cost ($\gamma_{\mathrm{v}}=10\%$). The figure shows that the energy consumption per bit of developing a 3 layer - 10 node MLP is $2.41\times 10^{-6}[J/b]$, which is approximately $3.7$ times higher than the energy consumption per bit of one inference ($E_{\mathrm{inf,p,b}}=6.57\times 10^{-7}[J/b]$). Furthermore, we can observe that the energy consumption per bit decreases with more inferences, dropping from $2.11\times 10^{-6}[J/b]$ for $100$ inferences to $7.73\times 10^{-7}[J/b]$ for $1000$ inferences, which is a $2.73$ times of improvement on energy efficiency. Moreover, we can observe that models with architectural blocks that require higher computational complexity need significantly more energy per bit during development compared to simpler models and their energy cost per bit of one inference. For instance, the transformer consumes $3.74\times 10^{-4}\leavevmode\nobreak\ [\mathrm{J/b}]$, which means it is nearly $165$ and $20.7$ times less efficient than the MLP in making one inference. This is also confirmed in Fig. 9. As the number of inferences $\gamma$ in the current model increases, the energy efficiency of the AIoT lifecycle improves.

### v3

@@ VI Energy Cost of Training
To train a model, training data and a machine learning techniques is needed. In this study, we consider neural network architectures, centralized offline training, however, extensions of this work could consider online, distributed, federated or transfer approaches as well. As discussed in Section V, the training data typically goes through a preprocessing phase and then split into training and evaluation sets, with a $\beta$ split ratio as:
$N_{\mathrm{S}}=\underbrace{\beta N_{\mathrm{S}}}_{N_{\mathrm{S,T}}}+\underbrace{(1-\beta)N_{\mathrm{S}}}_{N_{\mathrm{S,E}}},$ (10)
where $N_{\mathrm{S,T}}$ and $N_{\mathrm{S,E}}$ represent the number of samples for training and evaluation, respectively. While many neural network architectures are available for AI models development, in this work, we focus on four representative types. Namely, Multilayer Perceptrons [11], Convolutional Neural Networks [38], Kolmogorov–Arnold Networks [39], and transformers [40]. Training a neural network involves two fundamental phases: forward propagation and backward propagation, depicted with blue and red in Fig. 6, respectively, on an example MLP. During the forward propagation, input data passes through the network layers. Each layer performs specific tasks, such as computing a linear combination of weights and biases, followed by applying activation functions in dense (fully connected) layers. Convolutional layers perform convolutional operations, while pooling mechanisms are applied on pooling layers. At the end of this phase, the loss is calculated by comparing the predicted outputs to the actual target values. Subsequently, the model performs backward propagation, where it calculates gradients of the loss with respect to each parameter, updates these gradients in reverse through the layers, and adjusts the weights and biases accordingly. This process completes one epoch. Consequently, we can calculate the energy costs associated with forward and backward propagation, leading to an understanding of the overall energy consumption for training, model evaluation, and subsequent inference. In the following sections, we showcase the computational complexity of training and evaluating the aforementioned Neural Networks.
@@ VI-A Computational Complexity of MLP, CNN, KAN and transformer
$M_{\mathrm{MLP}}$ $=\sum_{l=1}^{L-1}\Bigg{(}\underbrace{(2M_{l-1}M_{l})}_{\text{Weight and bias}}+\underbrace{2M_{l}}_{\text{Summation and activation}}\Bigg{)},$ (11)
where the first term in the sum represents one multiplication and one addition ($=2$ FLOPs) related to the weight and bias corresponding to the edges times the number of nodes in that layer ($=M_{l-1}M_{l}$), while the second term represents the summation of all the values of the incoming edges and the application of the activation function ($=2$ FLOPs) times the number of nodes in that layer ($M_{l}$).
@@
$M_{\mathrm{MLP,FP}}$ $=N_{\mathrm{epochs}}N_{\mathrm{batch}}BM_{\mathrm{MLP}}$ (12)
$=N_{\mathrm{epochs}}N_{\mathrm{S,T}}M_{\mathrm{MLP}}.$
@@
$M_{\text{CONV}}=\underbrace{\Bigg{(}\frac{I_{\text{r}}-K_{\text{r}}+2P_{\text{r}}}{S_{\text{r}}}+1\Bigg{)}}_{\text{Output height}}\underbrace{\Bigg{(}\frac{I_{\text{c}}-K_{\text{c}}+2P_{\text{c}}}{S_{\text{c}}}+1\Bigg{)}}_{\text{Output width}}$ (13)
$\underbrace{\Bigg{(}C_{in}K_{\text{r}}K_{\text{c}}+1\Bigg{)}}_{\text{Computation per filter}}\underbrace{N_{\text{f}}}_{\text{No. of filters}},$
@@
$M_{\text{POOL}}=\underbrace{\Bigg{(}\frac{I_{\text{r}}-K_{\text{r}}}{S_{\text{r}}}+1\Bigg{)}}_{\text{Output height}}\underbrace{\Bigg{(}\frac{I_{\text{c}}-K_{\text{c}}}{S_{\text{c}}}+1\Bigg{)}}_{\text{Output width}}\underbrace{C_{\text{in}}}_{\text{Input channels}}.$ (14)
@@
$M_{\mathrm{CNN,FP}}$ $=N_{\mathrm{epochs}}N_{\mathrm{S,T}}$ (15)
@@
$M$ =${}_{\mathrm{KAN}}=\sum_{l=1}^{L-1}(\underbrace{M_{\mathrm{NLF}}M_{l-1}}_{\text{From B-spline}}$ (16)
$+\underbrace{\left(M_{l-1}M_{l}\right)}_{\text{Input and output}}\underbrace{\left[9K\left(G+1.5K\right)+2G-2.5K+3\right])}_{\text{B-spline andgrid}}$
$=\sum_{l=1}^{L-1}(M_{\mathrm{NLF}}M_{l-1}+\left(M_{l-1}M_{l}\right)M_{\mathrm{B}}),$
@@
$M_{\mathrm{KAN,FP}}=$ $N_{\mathrm{epochs}}N_{\mathrm{S,T}}M_{\text{KAN}}.$ (17)
@@
Based on the attention mechanism, the transformers [40] represent a paradigm shift in deep learning, replacing traditional architectures such as CNN with self-attention mechanisms. While enabling significant breakthroughs in AI through efficient parallel processing and robust modeling of long-range dependencies, they are relatively complex. In addition to attention layers, the transformer architecture also includes MLP, normalization, and positional encoding, making the analytic expression for the FLOP computation intractable. Following existing literature [43] and approximations  [^4] https://www.gaohongnan.com/playbook/training/how_to_calculate_flops_in_transformer_based_models.html, we approach the estimation as follows: $M_{\text{ATT}}=2(\underbrace{CN_{\text{embed}}3N_{\text{embed}}}_{\text{K, Q, V positional embedding}}+\underbrace{C^{2}N_{\text{embed}}}_{\text{Attention scores}}$ (18) $+\underbrace{N_{\text{head}}C^{2}N_{\text{embed}}/N_{\text{head}}}_{\text{reduce}}+\underbrace{CN_{\text{embed}}N_{\text{embed}}}_{\text{Projection}}),$ where $C$ is the context length, $N_{\text{embed}}$ is the size of the embedding and $N_{\text{head}}$ is the number of heads. The complexity of the decoder in a transformer is then given by:
$M_{\text{TR}}=N_{\text{decoder\_blocks}}(M_{\text{ATT}}+\underbrace{4CN_{\text{embed}}FFS}_{\text{MLP blocks}}),$ (19)
@@
$M_{\mathrm{TR,FP}}=$ $N_{\mathrm{epochs}}N_{\mathrm{S,T}}M_{\text{TR}}.$ (20)
@@ VI-B Model Training
To determine the computational complexity of training a model, backward propagation needs to also be considered. The standard approximation, also confirmed in [44], is that the backward propagation requires two times the floating-point operations per cycle per core (FLOPs) of the forward propagation. In particular, backward propagation requires matrix multiplications for updating the weights and propagating the gradient. Thus, we can approximate the computing complexity of training a model as:
@@ VI-C Model Evaluation
${E_{\mathrm{eval}}=\frac{M_{\mathrm{model}}N_{\mathrm{S,E}}}{PU_{\mathrm{performance}}}}.$ (23)
@@ VII Energy Cost of Inference
The complexity of making an inference depends on the number of FLOPs required for forward propagating with input sample size $N_{\mathrm{I,P}}$, i.e., $N_{\mathrm{inf}}=M_{\mathrm{model}}N_{\mathrm{I,P}}$. For example, when we adopt the aforementioned MLP model with $N_{\mathrm{I,P}}=51$, the complexity is calculated as $N_{\mathrm{inf}}=23651=12,036$ FLOPs. Therefore, the energy consumption of the forward propagation and the corresponding inference can be calculated as:
$E_{\mathrm{inf}}=\frac{M_{\mathrm{model}}N_{\mathrm{I,P}}}{PU_{\mathrm{performance}}}.$ (24)
We can then observe from Eq. 23 and 24 that the cost of inference is the same as in evaluation when $N_{\mathrm{S,E}}$ is equal to $N_{\mathrm{I,P}}$. Thus, we can conclude that the computational complexity between finishing one training and inference is on the magnitude of $3N_{\mathrm{epochs}}N_{\mathrm{S,T}}/N_{\mathrm{I,P}}$. This indicates that the training process involves significantly more computational operations, particularly when the number of epochs or $N_{\mathrm{S,T}}>>N_{\mathrm{I,P}}$. On the other hand, the energy consumption per bit depends on the model complexity and hardware capability. Moreover, the factor of three in the energy consumption per bit during training versus inference highlights the additional computational load inherent to training.
VIII eCAL: the Energy Cost of AI Lifecycle
@@
In this section, we describe the end-to-end energy consumption of the AI model lifecycle for making inferences with the aforementioned neural network architectures. First, we consider the energy consumption involved in developing the model, as illustrated on the left in Fig. 1, which can be expressed as follows:
@@
To compare the energy consumption of all the data manipulation components required for developing and deploying an AI-enhanced system, we assume that the models are trained and evaluated using 5-fold cross-validation (as discussed in Section VI). The models require $256$ samples with an input size of $10$, i.e., $N_{\mathrm{S}}I_{\mathrm{S}}=2560$, collected from a device over a wireless network configured with the following parameters: $P_{\mathrm{T}}=200\leavevmode\nobreak\ [mW]$, $P_{\mathrm{R}}=5\leavevmode\nobreak\ [mW]$, $R_{\mathrm{T}}=R_{\mathrm{R}}=10\leavevmode\nobreak\ [Mbps]$, $N_{\mathrm{\_,cycle},l}=100\leavevmode\nobreak\ [c/b]$, and $P_{\mathrm{\_,cycle}}=10^{-10}\leavevmode\nobreak\ [W/c]$. We also assume the link between the device and the server is stable, and there is no retransmission needed. Each layer introduces DP and CP overhead of $10\%$ and $5\%$, respectively. Furthermore, normalization is utilized as the data transformation procedure, as discussed in Section V. Finally, model training requires $10$ epochs ($N_{\mathrm{epochs}}=10$) and utilizes $5$-fold validation. The energy consumption of this example is shown in Table III and reveals the contribution of each component to the total energy cost of developing models. Specifically, data collection dominates the total energy cost for developing a very simple model based on a 3-layer, 10-nodes-per-layer MLP architecture. For the same parameters, the relative contribution of the data collection decreases for the KAN, CNN and transformer architectures, making the energy cost of training the primary contributor. The eCAL calculator enables changing the hyperparameters of the four architectures while also adding well known CNN architectures such as ResNets [45], VGGs [46] or transformer-based such as Baichuan 2 [47].
Once the model is developed, it is packaged and deployed in the communication system, where it can produce inference in the form of continuous or discrete outputs. During its operational lifecycle, the model will be presented with some input data and requested to produce the corresponding inference. Thus, we can express the associated energy cost as follows:
@@
This means that we consider the energy cost of inference before the current model needs to be updated. The corresponding energy per bit of inference, once deployed, is calculated as follows:
@@
Next, we calculate the total energy cost over the lifetime of an AI model in the system, also referred to as $eCAL_{\mathrm{abs}}$ (measured in [$J$]). It is defined as the sum of energy consumed during the model development, $E_{\mathrm{D}}$ and the energy required for performing an inference $E_{\mathrm{inf,p}}$, multiplied by the number of times $\gamma$ the inference is performed with the currently deployed model. In addition, as edge cloud systems may use virtualization in developing and operating the models, we enable considering the impact of such technologies through a factor. As a result, $eCAL_{\mathrm{abs}}$ can be expressed as:
@@
Figure 8: Comparison of energy consumption per bit in different data manipulation components over the lifecycle of the AI model.
Figure 9: Energy cost of AI model lifecycle ($eCAL$) for different models over the number of inferences ($\gamma$), log scale on both axes.
Finally, the proposed metric computing the corresponding energy consumption per bit over the lifecycle of the currently deployed AI model can be expressed as:
@@
To further quantify this relationship, we first compare the outcomes of Eq. 26, Eq. 28, and Eq. 30, depicted in Fig. 8, where we assume the virtualization is contributing additional $10\%$ to the total energy cost ($\gamma_{\mathrm{v}}=10\%$). The figure shows that the energy consumption per bit of developing a 3 layer - 10 node MLP is $2.41\times 10^{-6}[J/b]$, which is approximately $3.7$ times higher than the energy consumption per bit of one inference ($E_{\mathrm{inf,p,b}}=6.57\times 10^{-7}[J/b]$). Furthermore, we can observe that the energy consumption per bit decreases with more inferences, dropping from $2.11\times 10^{-6}[J/b]$ for $100$ inferences to $7.73\times 10^{-7}[J/b]$ for $1000$ inferences, which is a $2.73$ times of improvement on energy efficiency. Moreover, we can observe that models with architectural blocks that require higher computational complexity need significantly more energy per bit during development compared to simpler models and their energy cost per bit of one inference. For instance, the transformer consumes $3.74\times 10^{-4}\leavevmode\nobreak\ [\mathrm{J/b}]$, which means it is nearly $165$ and $20.7$ times less efficient than the MLP in making one inference. This is also confirmed in Fig. 9. As the number of inferences $\gamma$ in the current model increases, the energy efficiency of the AI model lifecycle improves.